# High-Performance Wavelength Division Multiplexers Enabled by Co-Optimized Inverse Design

### v1

@@ 5.1 Device Fabrication
The designed devices are fabricated on a silicon on insulator (SOI) substrate with a silicon device layer of 220 nm. The inverse-designed WDM and Bragg structures are patterned using electron beam lithography and positive resist (ZEP520A). The resist is developed in solvents (xylene then methyl isobutyl ketone (MIBK) and isopropyl alcohol (IPA)) and the silicon is etched using reactive ion etching (RIE). A HBr and Cl2 chemistry is used for the RIE. The resist is removed (using Remover 1165) and the substrate is cleaned in piranha solution (sulfuric acid and hydrogen peroxide) and an HF dip prior to oxide cladding. The silicon structures are cladded in oxide through spin coating of hydrogen silsesquioxane (HSQ) and annealing in nitrogen gas at 900∘ C.
@@ 6.2 Author Contributions
S.M., G.A., and J.V. conceived of the project. S.M. designed, simulated, and measured the devices. S.M. and S.E. fabricated the devices. J.G. designed and fabricated the frequency comb source, and J.G. and S.M. conducted the frequency comb measurements. G.A. and J.V. supervised the project. All authors contributed to the writing of the manuscript and the analysis of the results.

### v2

@@ 4.1 Device Fabrication
The silicon devices are fabricated on a silicon on insulator (SOI) substrate with a silicon device layer of 220 nm. The inverse-designed WDM and Bragg structures are patterned using electron beam lithography and positive resist (ZEP520A). The resist is developed in solvents (xylene then methyl isobutyl ketone (MIBK) and isopropyl alcohol (IPA)) and the silicon is etched using reactive ion etching (RIE). A HBr and Cl2 chemistry is used for the RIE. The resist is removed (using Remover 1165) and the substrate is cleaned in piranha solution (sulfuric acid and hydrogen peroxide) and an HF dip prior to oxide cladding. The silicon structures are cladded in oxide through spin coating of hydrogen silsesquioxane (HSQ) and annealing in nitrogen gas at 900∘C.
The silicon nitride devices are fabricated on a LPCVD silicon nitride layer that is deposited on a silicon substrate with 5 m SiO2 layer on top. The silicon nitride is deposited at 830∘C furnace yielding low-loss stoichiometric silicon nitride. As with the silicon devices, the WDMs are patterned and developed using the same electron beam lithography process described above. The silicon nitride is etched using reactive ion etching with a CF4 and CHF3 chemistry. The resist is removed (using Remover 1165) and the substrate is cleaned in piranha solution (sulfuric acid and hydrogen peroxide) and an HF dip.
@@ 4.2 Experimental Setup
5 Data Availability
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
6 Code Availability
The code generated for the current study are available from the corresponding author on reasonable request.
@@ 7.2 Author Contributions
S.M., G.A., and J.V. conceived of the project. S.M. designed, simulated, and measured the devices. S.M., J.G., and S.E. fabricated the devices. J.G. designed and fabricated the frequency comb source, and J.G. and S.M. conducted the frequency comb measurements. G.A. and J.V. supervised the project. All authors contributed to the writing of the manuscript and the analysis of the results.
7.3 Competing Interests
The authors declare the following competing interest:
Patent Application No. PCT/US2025/036120.